OpenStudy (anonymous):

I'm trying to embed a whole bunch of equations into MS word 2010. Are there any tips to make them look as nicely as they would if I used \( \LaTeX{} \)?

OpenStudy (anonymous):

don't use MS word

OpenStudy (anonymous):

In an ideal world, I would go with pdflatex + a text editor, but my collaborators require me to use MS Word 2010 :) How would I produce beautiful documents now?

OpenStudy (anonymous):

For example, how would I place \[\begin{align*} \frac{n\sum{x^2}-(\sum{x})^2}{n(n-1)} \end{align*}\] in word?